### 01 Apr 2026 v1

@@ Abstract
Figure 1: Neural Harmonic Textures for novel view synthesis. We attach learnable feature vectors (right) to the virtual vertices of bounding tetrahedra encapsulating each primitive (center). After harmonic encoding and accumulation along the ray, a small neural network decodes the resulting signal into RGB color in a deferred manner (left).
@@ 5.1.1 Optimization
where $P$ is the number of primitives and $\mathbf{s}_{i}$ is the scale vector of primitive $i$. Our final loss function is:
@@ 5.1.2 Implementation details
We implement our method in gsplat [ye2025gsplat], using the 3DGUT [3dgut] formulation, to which we add custom CUDA kernels for the forward and backward logic described in Sec.˜4. To reduce register pressure, we use half-precision memory fetches on feature vectors during rasterization, for both forward and backward kernels. Our neural network uses tiny-cuda-nn’s [mueller2022instant] JIT-compiled cooperative vector MLPs, trained and evaluated in half precision (FP16). We apply an automatic scaling factor to the loss values in order to improve half-precision training stability [micikevicius2017mixed].
@@ 5.1.3 Results
We present the results of our method against a number of previous works using pure neural fields [barron2022mipnerf360], hashgrid-encoded neural fields [mueller2022instant, zipnerf], pure primitive-based methods [2dgs, 3dgsmcmc, 3dgut, betasplatting, trianglesplatting] with regular Spherical Harmonics (SH) appearance or more complex appearance models [sphericalVoronoi, texturedgaussians], and closer to us, methods mixing primitives (e.g. for acceleration) and neural fields [nestsplatting, radiancemesh] (Tab.˜1). We consistently outperform all previous works in a varied set of real datasets [barron2022mipnerf360, Knapitsch2017, hedman2018deep]. We include a selection of views in Fig.˜7. Our method particularly excels at high-frequency detail modelling, capturing specular highlights and reflections with superior fidelity.
Table 1: Comparison on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep], on Neural Field-style methods, primitive-based methods, and mixed neural field/primitive-based methods. For our method, we use 64 features per primitive (16 per vertex), with a 128-wide $\times$ 3 hidden layers MLP, which results in a roughly similar number of total paramaters as previous primitive-based approaches on average. For indoor scenes, which typically span a smaller spatial volume, we use 2M primitives, while for larger outdoor scenes we increase the primitive count to 4.5M. To regularize training in high-parallax regimes (e.g., indoor datasets), we encode the central camera ray rather than the individual ray direction. Finally, to ensure a better coverage in terms of training epochs, we train indoor datasets longer (45k iterations for $\sim 290$ images) than outdoor (20k for $\sim 175$ images), following Spherical Voronoi [sphericalVoronoi]. Note that convergence of an MLP-based appearance model can be slower than SH-based ones.
@@
2DGS [2dgs] 27.48 0.816 0.263 22.85 0.827 0.244 29.56 0.904 0.325
3DGS-MCMC [3dgsmcmc] 28.21 0.841 0.214 24.46 0.866 0.174 29.49 0.912 0.306
3DGUT-MCMC [3dgsmcmc] 28.08 0.837 0.218 24.20 0.861 0.180 29.87 0.913 0.309
@@
Neural Harmonic Textures (Ours) 29.01 0.840 0.212 25.68 0.882 0.141 30.94 0.919 0.302
@@
Per-scene breakdowns and more details are provided in the appendix (Appendix Tab.˜10 and Tab.˜11).
@@ 5.1.4 Ablations
Appendix Tab.˜14 isolates the individual contributions of the training strategies detailed in Sec.˜5.1.2: learning rate scheduling, EMA on the MLP weights, loss scaling, the color-refinement phase, and opacity and scale regularization.
@@
Tab.˜19 and Tab.˜19 in the appendix explore the impact of varying the per-primitive feature dimension and the MLP architecture, respectively, to illustrate the resulting quality-speed trade-offs.
@@ 5.2.2 Semantic Field Reconstruction
We thus test our method on joint novel-view synthesis of RGB radiance and LSEG [LSEG] 512-wide semantic features. Due to the limited resolution of LSEG feature maps, we render and predict at the RGB resolution, but bilinearly downsample the spatial dimension of the resulting feature maps in order to supervise on ground truth features. We compare against Feature 3DGS [feature3dgs], a hybrid method relying on 128-wide semantic feature vector per primitive plus a CNN to jointly downsample the spatial dimension and upsample the feature dimension of the resulting rasterized image up to the full 512 dimensions, and another set of 48 SH basis weights for regular RGB color. Results for the aggregated MipNeRF360 dataset can be seen in Tab.˜6. We afford Feature-3DGS twice our training time budget, up to 7K iterations, slightly more than in the original paper (5K), to accommodate for the larger-scale datasets. We further ablate the per-primitive feature dimensionality for this task in Appendix Tab.˜12. Our results show substantial improvement over Feature 3DGS, and the capability to easily scale upwards or downwards in performance or memory, in a completely detached manner from the nature of the data being encoded. It is particularly interesting to note how small the drop in RGB quality is, which points to either high correlation among feature vectors or excessive free bandwidth in our RGB experiments. Our features can then be used in downstream tasks such as scene editing or semantic segmentation.
@@ Appendix 0.A Additional Results
In Tab.˜10 we provide a detailed breakdown of Tab.˜2 from the main paper, with per-scene results on MipNeRF360, Tanks and Temples and Deep Blending. For this comparison, we isolate completely the effect of our method by ensuring an even setup for all methods. All four approaches are implemented under the same framework (gsplat), trained for 30k iterations, using 1M primitives for all scenes, sharing training parameters among all scenes, and limiting to 48 features per primitive reserved for appearance. All results presented in this paper compute the LPIPs metric using the VGG model with normalized inputs to the range of [1,1][-1,1]. Furthermore, in Tab.˜8 we provide another standardized comparison on all three benchmarks, only this time we match the same primitive count in each scene as the high-quality reported baselines in their respective papers, training for the same amount of iterations. We include our method with both 48 and 64 features per primitive (48F, 64F).
@@
Tab.˜11 details the per-scene results of main Tab.˜1, which uses our split indoor/outdoor training configuration. This is our higher quality setup, which adapts to the different scene scale and dataset sizes of outdoor and indoor datasets.
@@
Table 11: Per-scene results of NHT (Ours) with split training configurations on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep]. We adapt primitive count and training length per dataset to accommodate different dataset sizes and spatial scales, but keep the same learning rates and other hyperparameters. Indoor datasets benefit from encoding the central camera ray, instead of per-ray directions, due to their high parallax, acting as a regularization on ray orientation.
@@
M360 Outdoor (4.5M, 20k iters, per-ray)
garden 28.33 0.881 0.106
bicycle 25.88 0.788 0.216
stump 27.06 0.795 0.223
treehill 23.36 0.670 0.306
flowers 22.08 0.640 0.330
@@
bonsai 35.73 0.964 0.192
counter 31.00 0.934 0.193
kitchen 33.59 0.945 0.123
room 34.03 0.947 0.221
@@
M360 Outdoor Avg 25.34 0.755 0.236
M360 Indoor Avg 33.59 0.947 0.183
M360 Total 29.01 0.840 0.212
@@ 0.A.1 Semantic Reconstruction
where $\hat{f}$ and $f^{*}$ are predicted and ground-truth LSEG feature maps, and coscos their cosine similarity. We compute these losses at the native LSEG resolution: the model renders at RGB resolution (high); then, we bilinearly downsample the predicted feature map to match the (lower-resolution) LSEG targets. Tab.˜13 lists the LSEG-specific hyperparameters used in our final training. We largely kept the same configuration from our pure RGB radiance field experiments; further tuning for this specific use case may render higher performance.
@@
Finally, in Tab.˜12 we ablate the feature dimensionality for the joint RGB and LSEG semantic reconstruction task.
@@ Appendix 0.B Feature interpolation
where $A$ is the total area of the triangle and $A_{i}$ is the area of the sub-triangle formed by the evaluation point $\mathbf{p}$ and the edge opposite to vertex $i$. For a triangle with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$:
@@
where $V$ is the total volume of the tetrahedron and $V_{i}$ is the volume of the sub-tetrahedron formed by the evaluation point $\mathbf{p}$ and the face opposite to vertex $i$. For a tetrahedron with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3},\mathbf{v}_{4}$:
@@ 0.B.1 Optimization choices and hyperparameters
Tab.˜14 ablates the individual contributions of the training strategies described in the main paper (learning rate scheduling, EMA on the MLP weights, loss scaling, color-refinement phase, and opacity and scale regularization). Tab.˜15 lists the full set of hyperparameters and optimization details used for radiance field reconstruction.
@@
SSIM weight $\lambda$ (D-SSIM vs $\mathrm{L}_{1}$) 0.2
@@
Scale regularization $\lambda_{s}$ 0.01
@@
Scales $4\times 10^{-3}$
Opacities $4\times 10^{-2}$
Quaternions $8\times 10^{-4}$
Deferred features $1.9\times 10^{-2}$
Deferred MLP $7.2\times 10^{-4}$
@@ 0.C.5 Post-Training Compression
We compress vertex positions to UInt16, vertex features to uniform Int8 (with a per-channel scale and offset), and MLP weights to FP16. An entropy coder (Zstandard, level 19) is applied to the serialized payload. Combining these techniques, the resulting images can be compressed by an additional factor of approximately $3\times$ beyond the training result, with negligible quality loss. Tab.˜20 summarizes the results of the post-training compression techniques.
@@ 0.C.6 Per-Image Results
Tabs.˜21 and 22 provide per-image results for the 2D image fitting experiment at $10\times$ and $100\times$ compression ratios, respectively. We provide visual comparisons in Fig.˜8 and Fig.˜10. Our method achieves competitive PSNR in both $\mu$-law and tonemapped spaces while substantially outperforming Instant NGP in perceptual quality (LPIPS). The advantage is most pronounced at high compression ratios, where the non-overlapping mesh topology and $C^{1}$ feature interpolation avoid the ringing and block artifacts that plague hash-grid methods.
@@
Tab.˜23 provides the full set of hyperparameters for the 2D image fitting experiment.

### 08 Apr 2026 v2

@@ Abstract
Figure 1: Neural Harmonic Textures for novel view synthesis. We attach learnable feature vectors (right) to the virtual vertices of bounding tetrahedra encapsulating each primitive (center). After harmonic encoding and accumulation along the ray, a small neural network decodes the resulting signal into RGB color in a deferred manner (left). Source code and further results are available at https://research.nvidia.com/labs/sil/projects/neural-harmonic-textures/.
@@ 5.1.1 Optimization
where $P$ is the number of primitives and $\mathbf{s}_{i}$ is the scale vector of primitive ii. Our final loss function is:
@@ 5.1.2 Implementation details
We implement our method in gsplat [ye2025gsplat], using the 3DGUT [3dgut] formulation, to which we add custom CUDA kernels for the forward and backward logic described in Sec.˜4. To reduce register pressure, we use half-precision memory fetches on feature vectors during rasterization, for both forward and backward kernels. Our neural network uses tiny-cuda-$n$’s [mueller2022instant] JIT-compiled cooperative vector MLPs, trained and evaluated in half precision (FP16). We apply an automatic scaling factor to the loss values in order to improve half-precision training stability [micikevicius2017mixed].
@@ 5.1.3 Results
We present the results of our method against a number of previous works using pure neural fields [barron2022mipnerf360], hashgrid-encoded neural fields [mueller2022instant, zipnerf], pure primitive-based methods [2dgs, 3dgsmcmc, 3dgut, betasplatting, trianglesplatting] with regular Spherical Harmonics (SH) appearance or more complex appearance models [sphericalVoronoi, texturedgaussians], and closer to us, methods mixing primitives (e.g. for acceleration) and neural fields [nestsplatting, radiancemesh] (Tab.˜1). In this table, we use the originally provided JPEG images for proper comparison with previous works in MipNeRF360; every other result presented in this section, unless otherwise stated, uses gsplat’s default downscaling for all methods tested. We measure the performance delta of this choice in the supplementary (Tab.˜10). We consistently outperform all previous works in a varied set of real datasets [barron2022mipnerf360, Knapitsch2017, hedman2018deep]. We include a selection of views in Fig.˜7. Our method particularly excels at high-frequency detail modelling, capturing specular highlights and reflections with superior fidelity.
Table 1: Comparison on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep], on Neural Field-style methods, primitive-based methods, and mixed neural field/primitive-based methods. For the MipNeRF360 datasets, we disable gsplat’s default downscaling and instead train and evaluate directly on the provided JPEG-compressed reference images following prior work. We measure the effect in Tab.˜10. For our method, we use 64 features per primitive (16 per vertex), with a 128-wide $\times$ 3 hidden layers MLP, which results in a roughly similar number of total parameters as previous primitive-based approaches on average. For indoor scenes, which typically span a smaller spatial volume, we use 2M primitives, while for larger outdoor scenes we increase the primitive count to 5M. To regularize training in high-parallax regimes (e.g., indoor datasets), we encode the central camera ray rather than the individual ray direction. Finally, to ensure similar training epochs, we train indoor datasets longer (45k iterations for $\sim 290$ images) than outdoor (25k for $\sim 175$ images), following Spherical Voronoi [sphericalVoronoi]. Unlike Spherical Voronoi, we do not fine tune learning rates per dataset. Note that convergence of an MLP-based appearance model can be slower than SH-based ones.
@@
2DGS [2dgs] 27.22 0.804 0.275 22.85 0.827 0.244 29.56 0.904 0.325
3DGS-MCMC [3dgsmcmc] 27.99 0.830 0.229 24.46 0.866 0.174 29.49 0.912 0.306
3DGUT-MCMC [3dgsmcmc] 27.82 0.826 0.233 24.20 0.861 0.180 29.87 0.913 0.309
@@
Neural Harmonic Textures (Ours) 28.74 0.834 0.216 25.68 0.882 0.141 30.94 0.919 0.302
@@
Per-scene breakdowns and more details are provided in the appendix (Appendix Tab.˜11 and Tab.˜12).
@@ 5.1.4 Ablations
Appendix Tab.˜15 isolates the individual contributions of the training strategies detailed in Sec.˜5.1.2: learning rate scheduling, EMA on the MLP weights, loss scaling, the color-refinement phase, and opacity and scale regularization.
@@
Tab.˜20 and Tab.˜20 in the appendix explore the impact of varying the per-primitive feature dimension and the MLP architecture, respectively, to illustrate the resulting quality-speed trade-offs.
@@ 5.2.2 Semantic Field Reconstruction
We thus test our method on joint novel-view synthesis of RGB radiance and LSEG [LSEG] 512-wide semantic features. Due to the limited resolution of LSEG feature maps, we render and predict at the RGB resolution, but bilinearly downsample the spatial dimension of the resulting feature maps in order to supervise on ground truth features. We compare against Feature 3DGS [feature3dgs], a hybrid method relying on 128-wide semantic feature vector per primitive plus a CNN to jointly downsample the spatial dimension and upsample the feature dimension of the resulting rasterized image up to the full 512 dimensions, and another set of 48 SH basis weights for regular RGB color. Results for the aggregated MipNeRF360 dataset can be seen in Tab.˜6. We afford Feature-3DGS twice our training time budget, up to 7K iterations, slightly more than in the original paper (5K), to accommodate for the larger-scale datasets. We further ablate the per-primitive feature dimensionality for this task in Appendix Tab.˜13. Our results show substantial improvement over Feature 3DGS, and the capability to easily scale upwards or downwards in performance or memory, in a completely detached manner from the nature of the data being encoded. It is particularly interesting to note how small the drop in RGB quality is, which points to either high correlation among feature vectors or excessive free bandwidth in our RGB experiments. Our features can then be used in downstream tasks such as scene editing or semantic segmentation.
@@ Appendix 0.A Additional Results
In Tab.˜11 we provide a detailed breakdown of Tab.˜2 from the main paper, with per-scene results on MipNeRF360, Tanks and Temples and Deep Blending. For this comparison, we isolate completely the effect of our method by ensuring an even setup for all methods. All four approaches are implemented under the same framework (gsplat), trained for 30k iterations, using 1M primitives for all scenes, sharing training parameters among all scenes, and limiting to 48 features per primitive reserved for appearance. All results presented in this paper compute the LPIPs metric using the VGG model with normalized inputs to the range of [1,1][-1,1]. Furthermore, in Tab.˜8 we provide another standardized comparison on all three benchmarks, only this time we match the same primitive count in each scene as the high-quality reported baselines in their respective papers, training for the same amount of iterations. We include our method with both 48 and 64 features per primitive (48F, 64F).
@@
Tab.˜12 details the per-scene results of main Tab.˜1, which uses our split indoor/outdoor training configuration. This is our higher quality setup, which adapts to the different scene scale and dataset sizes of outdoor and indoor datasets.
Effect of reference image format.
As noted in Tab.˜1, MipNeRF360 images are distributed as pre-downscaled, re-compressed JPEG files. The default gsplat pipeline downscales and converts these to lossless PNGs before training and evaluation, which removes JPEG compression artifacts from the ground truth. In Tab.˜10 we compare NHT results under both protocols. Training and evaluating directly on the original JPEG references yields slightly lower PSNR (-0.28 dB), SSIM (-0.011), and higher LPIPS (++0.014) on average, as JPEG compression artifacts in the ground truth lower the effective quality ceiling.
Table 10: Effect of reference image format on MipNeRF360 evaluation for NHT (Ours). Downscaled PNGs uses gsplat’s default strategy, which downscales and converts the provided JPEG references to lossless PNGs before training and evaluation. JPEG (original) trains and evaluates directly on the provided JPEG-compressed references without any preprocessing. JPEG compression artifacts in the ground truth lower the effective quality ceiling, slightly degrading all three metrics.
[…]
Downscaled PNGs 29.022 0.8447 0.2026
JPEG (original) 28.738 0.8341 0.2164
$\Delta$ -0.284 -0.0106 ++0.0138
@@
Table 12: Per-scene results of NHT (Ours) with split training configurations on MipNeRF360 [barron2022mipnerf360], Tanks & Temples [Knapitsch2017], and Deep Blending [hedman2018deep]. We adapt primitive count and training length per dataset to accommodate different dataset sizes and spatial scales, but keep the same learning rates and other hyperparameters. Indoor datasets benefit from encoding the central camera ray, instead of per-ray directions, due to their high parallax, acting as a regularization on ray orientation. Trained and evaluated on the original JPEG-compressed reference images.
@@
M360 Outdoor (5M, 25k iters, per-ray)
garden 28.06 0.875 0.108
bicycle 25.53 0.782 0.203
stump 27.10 0.797 0.206
treehill 23.09 0.652 0.298
flowers 22.08 0.642 0.304
@@
bonsai 35.09 0.957 0.222
counter 30.83 0.925 0.220
kitchen 33.40 0.940 0.138
room 33.43 0.936 0.250
@@
M360 Outdoor Avg 25.17 0.749 0.224
M360 Indoor Avg 33.19 0.940 0.207
M360 Total 28.74 0.834 0.216
@@ 0.A.1 Semantic Reconstruction
where $\hat{f}$ and $f^{*}$ are predicted and ground-truth LSEG feature maps, and coscos their cosine similarity. We compute these losses at the native LSEG resolution: the model renders at RGB resolution (high); then, we bilinearly downsample the predicted feature map to match the (lower-resolution) LSEG targets. Tab.˜14 lists the LSEG-specific hyperparameters used in our final training. We largely kept the same configuration from our pure RGB radiance field experiments; further tuning for this specific use case may render higher performance.
@@
Finally, in Tab.˜13 we ablate the feature dimensionality for the joint RGB and LSEG semantic reconstruction task.
@@ Appendix 0.B Feature interpolation
where AA is the total area of the triangle and $A_{i}$ is the area of the sub-triangle formed by the evaluation point $\mathbf{p}$ and the edge opposite to vertex ii. For a triangle with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$:
@@
where $V$ is the total volume of the tetrahedron and $V_{i}$ is the volume of the sub-tetrahedron formed by the evaluation point $\mathbf{p}$ and the face opposite to vertex ii. For a tetrahedron with vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3},\mathbf{v}_{4}$:
@@ 0.B.1 Optimization choices and hyperparameters
Tab.˜15 ablates the individual contributions of the training strategies described in the main paper (learning rate scheduling, EMA on the MLP weights, loss scaling, color-refinement phase, and opacity and scale regularization). Tab.˜16 lists the full set of hyperparameters and optimization details used for radiance field reconstruction.
@@
SSIM weight $\lambda$ (D-SSIM vs $\mathrm{L}_{1}$) 0.1
@@
Scale regularization $\lambda_{s}$ 0.005
@@
Scales $5\times 10^{-3}$
Opacities $5\times 10^{-2}$
Quaternions $1\times 10^{-3}$
Deferred features $1.5\times 10^{-2}$
Deferred MLP $6.8\times 10^{-4}$
@@ 0.C.5 Post-Training Compression
We compress vertex positions to UInt16, vertex features to uniform Int8 (with a per-channel scale and offset), and MLP weights to FP16. An entropy coder (Zstandard, level 19) is applied to the serialized payload. Combining these techniques, the resulting images can be compressed by an additional factor of approximately $3\times$ beyond the training result, with negligible quality loss. Tab.˜21 summarizes the results of the post-training compression techniques.
@@ 0.C.6 Per-Image Results
Tabs.˜22 and 23 provide per-image results for the 2D image fitting experiment at $10\times$ and $100\times$ compression ratios, respectively. We provide visual comparisons in Fig.˜8 and Fig.˜10. Our method achieves competitive PSNR in both $\mu$-law and tonemapped spaces while substantially outperforming Instant NGP in perceptual quality (LPIPS). The advantage is most pronounced at high compression ratios, where the non-overlapping mesh topology and $C^{1}$ feature interpolation avoid the ringing and block artifacts that plague hash-grid methods.
@@
Tab.˜24 provides the full set of hyperparameters for the 2D image fitting experiment.